Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Delta course probabilies question . alternate solution help

Expert replies
by mrohekar » Sun May 06, 2007 8:08 am
I have arrived at a seperate way of arriving at the answer for this probability question.

I am terrible at probability so I am not quite sure if it is a fluke or my logic.

Here's the question:

The Delta Course
Daily GMAT Practice Question
May 5, 2007

The Full House Casino is running a new promotion. Each
person visiting the casino has the opportunity to play the
Trip Aces game. In Trip Aces, a player is randomly dealt
three cards, without replacement, from a deck of 8 cards. If
a player receives 3 aces, they will receive a free trip to
one of 10 vacation destinations. If the deck of 8 cards
contains 3 aces, what is the probability that a player will
win a trip?

A. 1/336
B. 1/120
C. 1/56
D. 1/720
E. 1/1440

The solution they gave is as follows:

The probability of an event A occurring is the number of
outcomes that result in A divided by the total number of
possible outcomes.

There is only one result that results in a win: receiving
three aces.

Since the order of arrangement does not matter, the number
of possible ways to receive 3 cards is a combination
problem.

The number of combinations of n objects taken r at a time is

C(n,r) = n!/(r!(n-r)!)

C(8,3) = 8!/(3!(8-3)!)
C(8,3) = 8!/(3!(5!))
C(8,3) = 40320/(6(120))
C(8,3) = 40320/720
C(8,3) = 56

The number of possible outcomes is 56.

Thus, the probability of being dealt 3 aces is 1/56.

The way I arrived at the solution is :

1) Probability of being dealt an ace out of 8 cards: 3/8
2) Probability of being dealth an ace out of 7 cards: 2/7
3) Probability of being dealt an ace out of 6 cards : 1/6
As cards are not being replaced.

Hence probability of being dealt 3 aces is: 3/8*2/7*1/6 = 1/56.

Your comments are appreciated.
Join the discussion
Source: — Problem Solving |

Actually both ways are good and correct.

If you do 8C3, you are essentially computing the probability over
all possible combinations of a set of 3 cards. There are a total of
8C3 combinations of which only 1 yields a trip-aces

Your method is equally correct where you are deciding on probability
one event at a time and using P(A)*P(B)*P(C) to compute the final
probability.

Use the one that you are most comfortable with
Join the discussion

by mrohekar » Sun May 06, 2007 8:23 am
Thanks for the quick reply...

I think I would always love to arrive at the answer with the least amount of calculations which I think mine does.

How does one make a judgement on finding the answer using the least amount of calculation ? Hope to find the answer before I give my GMAT :lol:
Join the discussion

by mendiratta » Tue May 08, 2007 9:31 pm
mrohekar wrote:Thanks for the quick reply...

I think I would always love to arrive at the answer with the least amount of calculations which I think mine does.

How does one make a judgement on finding the answer using the least amount of calculation ? Hope to find the answer before I give my GMAT :lol:
I too prefer your way for solving probability questions.
Join the discussion